OpenStudy (anonymous):

Which is the more reasonable measurement of the distance between the tracks on a CD: 1.6 ⋅ 10−3 mm or 1.6 ⋅ 103 mm?

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

1.6 . 10-3mm

OpenStudy (anonymous):

so it would be 1.6 ⋅ 10−3 mm

OpenStudy (perl):

10^3 mm = 10*10*10 mm = 1000 mm , which is pretty large distance between two tracks on a cd

OpenStudy (anonymous):

1.6 ⋅ 103 mm? would be more equivalent.

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@perl